Sara ordered 2 slices of pizza and a 12-ounce cola and paid $3.00. Sydney ordered 3 slices of pizza and 2 12-ounce colas for $4.75. How much does a slice of pizza cost?
1
Expert's answer
2012-11-07T09:49:00-0500

Let slice of pizza cost x and a 12-ounce cola cost y.
Then Sara's order cost: 2x+y=3
and Sydney's order cost: 3x+2y=4.75
From the first equation we have: y=3-2x and put this to the second equation:
3x+2(3-2x)=4.75
3x+6-4x=4.75
x=1.25
y=3-1.25*2=0.50
Answer: a slice of pizza cost $1.25 and a 12-ounce cola cost $0.50.